# An operator splitting analysis of Wasserstein–Fisher–Rao gradient flows

### v1

@@ 5 Convergence speed of Exact WFR: log-concave case
Figure 5.1: Comparison of exact KL decay (left) and symmetrised KL decay (right) for a 1D Gaussian with $m_{\pi}=20,C_{\pi}=100,m_{0}=0,C_{0}=1$. Left plot: We compare the exact decay of KL from Proposition 3.2 with the rates in (5.2) and (5.3) with $\delta=0.1,t_{0}=6.9$. Right plot: We compare the exact decay of the symmetrised KL obtained using the mean and covariance evolution in (3.1)–(3.2) with the rate in Proposition 5.1.
6 Splitting & convergence speed
Existing $\mathrm{KL}$ decay results for the W flow and FR flow developed in the literature cannot be straightforwardly utilised to characterise differences in convergence speed between the W-FR and FR-W scheme. In particular, combining existing exponential decay bounds for the KL divergence for the W part and the FR part would not allow us to isolate changes to convergence speed depending on the order as the exponents simply add (and commute). Hence, the first variation results developed in Proposition 2.2 and 2.3 will be critical in characterising how convergence speed is affected due to splitting, and the ordering of operators.
We utilise the variational formulae (2.9) and (2.14) to tighten the upper bound on the KL. We focus on one single split step of size $\gamma>0$ and consider the classical KL divergence $\mathrm{KL}(\mu||\pi)$, as a first step.
6.1 Convergence speed: W-FR
We first focus on the PDE (2.9) due to its easier form. It is possible to obtain a KL decay ODE by a straightforward calculation using (2.9), integration by parts, the sufficient tail decay of $\nu_{\gamma}$ and $\pi$ guaranteed by Assumption 1 and assuming sufficient regularity to interchange differentiation and integration.
Consider one step of the W-FR splitting (2.6) of size $\gamma>0$. Recall the split W-FR PDE in Proposition 2.2
$\partial_{\gamma}\nu_{\gamma}=f_{\textrm{W}}(\nu_{\gamma})+f_{\textrm{FR}}(\nu_{\gamma})+(e^{\gamma}-1)f_{\textrm{P}}(\nu_{\gamma}),$
where $f_{\textrm{P}}$ is in (2.12). Then, we have
$\frac{d}{d\gamma}\mathrm{KL}(\nu_{\gamma}||\pi)$ $=\int\log\frac{\nu_{\gamma}(x)}{\pi(x)}\left(f_{\textrm{W}}(\nu_{\gamma}(x))+f_{\textrm{FR}}(\nu_{\gamma}(x))+(e^{\gamma}-1)f_{\text{P}}(\nu_{\gamma}(x))\right)dx$ (6.1)
$=-\mathcal{I}(\nu_{\gamma}||\pi)-\text{Var}_{\nu_{\gamma}}\left[\log\frac{\nu_{\gamma}}{\pi}\right]+(e^{\gamma}-1)\text{Cov}_{\nu_{\gamma}}\left(\log\frac{\nu_{\gamma}}{\pi},\left|\nabla\log\frac{\nu_{\gamma}}{\pi}\right|^{2}\right),$
where $\mathcal{I}(\nu_{\gamma}||\pi)$ is the relative Fisher information.
It is clear that the behaviour of KL decay for the W-FR split will depend on the perturbation term $\text{Cov}_{\nu_{\gamma}}\left(g(\nu_{\gamma}),\left|\nabla g(\nu_{\gamma})\right|^{2}\right)$. This quantity measures the covariation between the log density ratio (when it is pointwise positive, it means that the current density is larger than the target density) and the squared score function (or magnitude of the score function), which is large in regions where the ratio $\frac{\nu_{\gamma}(x)}{\pi(x)}$ varies rapidly. In fact, if the log-ratio and the score are strongly negatively correlated, it is this negative correlation that contributes to $\text{Cov}_{\nu_{\gamma}}\left(g(\nu_{\gamma}),\left|\nabla g(\nu_{\gamma})\right|^{2}\right)$ being negative and driving the speed-up. In the below Lemma, we list conditions under which this covariance is negative. We emphasise that these conditions are sufficient, and establishing similar results for more relaxed conditions is left for future work. Given that a similar speed-up can be achieved for a 1D Gaussian mixture target (see Figure 1 (right) in [us_neurips]), we expect this to be possible.
However, we emphasise that negativity of the perturbation term is not sufficient to guarantee a speed-up, uniformly in time. A more precise characterisation, e.g. via comparison theorems, would require a lower-bound on the decay rate of the $\mathrm{KL}$. This aspect is left for future work.
Lemma 6.1.
The following results hold.
  1. (i)
    [us_neurips, Proposition 2] Consider $\nu(x)=\mathcal{N}(x;b,Q)$ and $\pi(x)=\mathcal{N}(x;m_{\pi},C_{\pi})$. If $Q\prec C_{\pi}$ then $\text{Cov}_{\nu}\left(g(\nu),\left|\nabla g(\nu)\right|^{2}\right)<0$.
  2. (ii)
    Denote $h(x):=-g(\nu(x))$. Assume $h$ is symmetric and $\nabla^{2}h\succeq\alpha I$. Then $\text{Cov}_{\nu}\left(g(\nu),\left|\nabla g(\nu)\right|^{2}\right)<0$.
@@ 6.1 Convergence speed: W-FR
We assume w.l.o.g. that $h$ is symmetric about 0 and that $h\geq 0$ since it is strongly convex and thus lower bounded. Since $h$ and $|\nabla h|^{2}$ are both decreasing in - and both increasing in + we can apply Chebyshev’s integral inequality (e.g. [egozcue2009some]) to each interval and obtain
$\int\nu(x)h(x)|\nabla h(x)|^{2}dx$ $=\frac{1}{\nu(_{-})}\int_{{}_{-}}\nu(x)h(x)|\nabla h(x)|^{2}dx+\frac{1}{\nu(_{+})}\int_{{}_{+}}\nu(x)h(x)|\nabla h(x)|^{2}dx$
$\geq\frac{1}{\nu(_{-})^{2}}\int_{{}_{-}}\nu(x)h(x)dx\int_{{}_{-}}\nu(x)|\nabla h(x)|^{2}dx$
$+\frac{1}{\nu(_{+})^{2}}\int_{{}_{+}}\nu(x)h(x)dx\int_{{}_{+}}\nu(x)|\nabla h(x)|^{2}dx.$
Due to the symmetry of $h$ we have $\int_{{}_{+}}\nu(x)h(x)dx=\int_{{}_{-}}\nu(x)h(x)dx$ and similarly for $|\nabla h|^{2}$. Using the fact that $\nu(_{+}),\nu(_{-})\leq 1$ since $\nu\in\mathcal{P}_{2}^{ac}(^{d})$, we finally have
$\int\nu(x)h(x)|\nabla h(x)|^{2}dx$ $\geq\int\nu(x)h(x)dx\int\nu(x)|\nabla h(x)|^{2}dx.$
Recalling that g=hg=-h, the result follows. ∎
Inequalities of the form $\text{Cov}_{\nu}(\alpha(X),\beta(X))<0$ for general measures $\mu$ and functions $\alpha,\beta$ are an active area of research [egozcue2011covariance, bonnefont2024covariance, egozcue2009some, saumard2014log] and we believe that bounds obtained from (6.1) can inform the choice of splitting to choose for a given pair of target $\pi$ and initial distribution $\mu_{0}$ provided these inequalities are obtained.
6.2 Convergence speed: FR-W
Similarly to the W-FR we can use the split PDE (2.14) to quantify the differen in KL decay due to splitting over one step of size $\gamma>0$.
Consider one step of the FR-W splitting (2.7) of size $\gamma>0$. We have
$\frac{d}{d\gamma}\mathrm{KL}(\eta_{\gamma}||\pi)$ $=\int\log\frac{\eta_{\gamma}(x)}{\pi(x)}\left(f_{\textrm{W}}(\eta_{\gamma}(x))+f_{\textrm{FR}}(\eta_{\gamma}(x))+\int_{0}^{\gamma}S_{\textrm{W}}\left(\gamma-\tau,-f_{\textrm{p}}(\eta_{\tau})\right)(x)d\tau\right)dx$
@@ 6.2 Convergence speed: FR-W
$=-\mathcal{I}(\eta_{\gamma}||\pi)-\text{Var}_{\eta_{\gamma}}\left[\log\frac{\eta_{\gamma}}{\pi}\right]-\int_{0}^{\gamma}\text{Cov}_{\eta_{\tau}}\left(S_{\textrm{W}}\left(\gamma-\tau,g(\eta_{\gamma})\right),|\nabla g(\eta_{\tau})|^{2}\right)d\tau.$
In the below Lemma, we list conditions under which
$\int_{0}^{\gamma}\text{Cov}_{\eta_{\tau}}\left(S_{\textrm{W}}\left(\gamma-\tau,g(\eta_{\gamma})\right),|\nabla g(\eta_{\tau})|^{2}\right)d\tau>0.$ (6.2)
Lemma 6.2.
Let $g(\eta_{\tau}(x))$ be symmetric with $\nabla^{2}g(\eta_{\tau})\succeq\alpha_{\tau}I$ for all $\tau\in[0,\gamma]$. Then (6.2) holds.

### v2

@@ 5.1 Exact WFR
Figure 5.1: Comparison of KL decay (left) and symmetrised KL decay (right) of the exact WFR for a 1D Gaussian with $m_{\pi}=20,C_{\pi}=100,m_{0}=0,C_{0}=1$. Left plot: Exact KL decay from Proposition 3.2 (blue) compared to rates in the literature, (5.2) (orange) and (5.3) (green) with $\delta=0.1,t_{0}=8.3$. Right plot: Exact symmetrised KL decay from (3.1)–(3.2) (blue) with our rate in Proposition 5.1 (orange).
5.2 W-FR split
We now utilise the variational formula (2.9) to obtain decay rates of symmetrised KL for the W-FR split, analogous to that of Proposition 5.1. Our result shows that under appropriate conditions, using $n$ steps of the W-FR split with time step $\gamma$ leads to a tighter upper bound on the decay of the symmetrised KL compared to the rate in Proposition 5.1. We note however that this is not necessarily a proof of speed-up. Such a characterisation, e.g. via comparison theorems, would require a lower-bound on the decay rate of the $\mathrm{KL}$ of both the exact and split flow. This aspect is left for future work.
From Proposition 2.2, it is clear that the behaviour of $J(\nu_{n},\pi)$ for the W-FR split scheme will depend on the perturbation term $f_{\text{P}}$. As will be seen later in this section, the analysis of this term will depend on the sign of $\text{Cov}_{\tilde{\nu}_{\tau}}\left(g(\nu_{1}),\left|\nabla g(\nu_{1})\right|^{2}\right)$ where $\tilde{\nu}_{\tau}\propto\nu_{1}^{(1-\tau)}\pi^{\tau}$ for all $\tau\in[0,1]$. This quantity measures the covariation between the log density ratio (when it is pointwise positive, it means $\nu_{1}$ is larger than the target density at that point) and the magnitude of the difference in score functions, which is large in regions where the ratio $\frac{\nu_{1}(x)}{\pi(x)}$ varies rapidly. When this covariance term is negative, regions where $\pi$ dominates $\nu_{1}$ tend to correspond to regions where local gradients of $\pi$ and $\nu_{1}$ differ, when measured against the interpolated density $\tilde{\nu}_{\tau}$.
Assumption 4.
Suppose $\mu_{0}$ and $\pi$ are chosen such that $\nu_{1}$, the solution of the W-FR split scheme after a single step of size $\gamma$, satisfies
$\text{Cov}_{\tilde{\nu}_{\tau}}\left(g(\nu_{1}),\left|\nabla g(\nu_{1})\right|^{2}\right)<0,$
where $g$ is as defined in (2.4) and $\tilde{\nu}_{\tau}\propto\nu_{1}^{1-\tau}\pi^{\tau}$ for all $\tau\in[0,1]$.
Deriving conditions on $\mu_{0}$ and $\pi$ such that Assumption 4 holds is a difficult task and is an active area of research [EGW+11, BHS24, EGW09, SW14]. Generally, inequalities of the form $\text{Cov}_{\nu}(\alpha(X),\beta(X))<0$ for general measures $\nu$ and functions $\alpha,\beta$ can be obtained under co-monotonicity assumptions on $\alpha,\beta$ [ARM93]. In Appendix D, we demonstrate it holds for $\nu(x)=\mathcal{N}(x;b,Q)$ and $\pi(x)=\mathcal{N}(x;m_{\pi},C_{\pi})$ with $0\prec Q\prec C_{\pi}$. Due to Proposition 3.2 ii), this property is maintained along both the exact and split flows for Gaussian $\mu_{0},\pi$, so long as it is satisfied initially. We speculate that similar results should hold when $g$ is strongly concave in $x$. A first step in this direction is examined in the 1D symmetric case in Appendix D.2. Establishing similar results for more relaxed conditions is left for future work. Given that a similar speed-up can be achieved for a 1D Gaussian mixture target (Figure 5.2), we expect this to be possible.
[…]
Figure 5.2: Ratio of KL from one step of W-FR scheme to KL from continuous time WFR as a function of step size $\gamma$ where $\pi$ is a 2 component univariate Gaussian mixture.
We now obtain the following result in direct analogy to Proposition 5.1. Assumption 4 ensures that for any $\gamma>0$, the decay rate of the upper bound on $J(\nu_{n},\pi)$ is higher than that of $J(\mu_{t},\pi)$ for any $t=n\gamma,n=1,2,3,\dots$
Proposition 5.2.
Assume the conditions of Corollary 4.2 and Assumption 4. Then the following decay result holds for $\nu_{n}$ the solution of the W-FR split scheme for all finite positive integer valued $n$,
$J(\nu_{n},\pi)<J(\mu_{0},\pi)e^{-n\gamma(\alpha_{\pi}+1)-c(e^{\gamma}-\gamma-1)}$
for some $c>0$ dependent on $\gamma$.
@@ 5.2 W-FR split
Given the assumptions, we may interchange differentiation and integration to obtain
$\frac{d}{d\gamma}J(\nu_{1},\pi)$ $=\int\left(\log\frac{\nu_{1}}{\pi}-\frac{\pi}{\nu_{1}}\right)(x)\left(f_{\textrm{W}}(\nu_{1}(x))+f_{\textrm{FR}}(\nu_{1}(x))+(e^{\gamma}-1)f_{\text{P}}(\nu_{1}(x))\right)dx$
$\leq-(\mathcal{I}(\nu_{1}||\pi)+\mathcal{I}(\pi||\nu_{1}))-J(\nu_{1},\pi)$
$+(e^{\gamma}-1)\left[\text{Cov}_{\nu_{1}}\left(\log\frac{\nu_{1}}{\pi},\left|\nabla\log\frac{\nu_{1}}{\pi}\right|^{2}\right)-\mathcal{I}(\pi||\nu_{1})+\mathcal{I}(\nu_{1}||\pi)\right].$
The term $\text{Cov}_{\nu_{1}}\left(\log\frac{\nu_{1}}{\pi},\left|\nabla\log\frac{\nu_{1}}{\pi}\right|^{2}\right)<0$ due to Assumption 4 with $\tau=0$. We will now show that $\mathcal{I}(\pi||\nu_{1})+\mathcal{I}(\nu_{1}||\pi)<0$ also under Assumption 4. Recall $\nu_{1}\propto e^{-V(x)}$ with $V(x)$ strongly convex due to Corollary 4.2. Define $U_{\tau}(x):=(1-\tau)V(x)+\tau V_{\pi}(x)$ and $\tilde{\nu}_{\tau}=\frac{e^{-U_{\tau}(x)}}{\int e^{-U_{\tau}(y)}dy}$. Clearly $\tilde{\nu}_{0}=\nu_{1}$ and $\tilde{\nu}_{1}=\pi$. Also define $F_{\tau}:=\int|\nabla g(\nu_{1}(x))|^{2}\tilde{\nu}_{\tau}(x)dx$ where $g(\nu_{1})=\log\frac{\nu_{1}}{\pi}$ and once again we have $F_{1}=\mathcal{I}(\pi||\nu_{1})$ and $F_{0}=\mathcal{I}(\nu_{1}||\pi)$. Note that since $\partial_{\tau}U_{\tau}=-V+V_{\pi}=g(\nu_{1})$,
$\partial_{\tau}\tilde{\nu}_{\tau}$ $=-\partial_{\tau}U_{\tau}\tilde{\nu}_{\tau}+\tilde{\nu}_{\tau}\int\partial_{\tau}U_{\tau}(x)\tilde{\nu}_{\tau}(x)dx=-g(\nu_{1})\tilde{\nu}_{\tau}+\tilde{\nu}_{\tau}\mathbb{E}_{\tilde{\nu}_{\tau}}\left[g(\nu_{1})\right]$
Then differentiating $F_{\tau}$ and swapping order of derivative and integral (permitted under the assumptions) yields
@@
$\frac{dF_{\tau}}{d\tau}=\int|\nabla g(\nu_{1}(x))|^{2}\partial_{\tau}\tilde{\nu}_{\tau}(x)dx=\int|\nabla g(\nu_{1}(x))|^{2}\left(-g(\nu_{1})\tilde{\nu}_{\tau}+\tilde{\nu}_{\tau}\mathbb{E}_{\tilde{\nu}_{\tau}}\left[g(\nu_{1})\right]\right)dx=-\text{Cov}_{\tilde{\nu}_{\tau}}\left(|\nabla g(\nu_{1})|^{2},g(\nu_{1})\right).$
Then under Assumption 4, we have that $\frac{dF_{\tau}}{d\tau}>0$ for all $\tau\in[0,1]$, then $F_{0}<F_{1}$ so that we immediately obtain $\mathcal{I}(\pi||\nu_{1})+\mathcal{I}(\nu_{1}||\pi)<0$. Inserting this back into the differential inequality yields
$\frac{d}{d\gamma}J(\nu_{1},\pi)$ $<-(\mathcal{I}(\nu_{1}||\pi)+\mathcal{I}(\pi||\nu_{1}))-J(\nu_{1},\pi)-(e^{\gamma}-1)\sigma_{1}(\gamma)$
for some $\sigma_{1}(\gamma)>0$ potentially dependent on $\gamma$. Integrating yields
$J(\nu_{1},\pi)$ $<J(\mu_{0},\pi)e^{-\gamma(\alpha_{\pi}+1)-\int_{0}^{\gamma}(e^{s}-1)\sigma_{1}(s)ds}$
where integral term is strictly negative. Applying an induction argument with Assumption 4 holding for $\mu_{0}=\nu_{i}$, $i=2,3,\dots$ yields the result with $c=\min_{n}\inf_{s}\sigma_{n}(s)>0$. ∎
Remark 5.3.
Assumption 4 can be weakened to require only $\text{Cov}_{\nu_{1}}\left(g(\nu_{1}),\left|\nabla g(\nu_{1})\right|^{2}\right)<0$ to obtain a result analogous to that of Proposition 5.2, but with restrictions on the size of $\gamma$.
Comparing the rate in Proposition 5.2 with that of the exact WFR flow in Proposition 5.1 we find that the evolution of the W-FR split allows us to get sharper rates. However, as discussed earlier, this does not guarantee faster convergence without a lower bound on $J(\nu_{n\gamma},\pi)$. It can however, inform the choice of initial distribution $\mu_{0}$. In fact, Assumption 4 cannot hold for the first splitting step if $\text{Cov}_{\mu_{0}}\left(g(\mu_{0}),\left|\nabla g(\mu_{0})\right|^{2}\right)>0$.
The analysis of the FR-W split is considerably more involved as the variational formula in Proposition 2.3 is less tractable. To control the perturbation terms arising from (2.12) one would in fact need to show that $\text{Cov}_{\tilde{\eta}_{\tau}}\left(g_{k}(\eta_{1}),g(\eta_{1})\right)<0$ for all $k\geq 1$, where $\tilde{\eta}_{\tau}\propto\eta_{1}^{(1-\tau)}\pi^{\tau}$ for all $\tau\in[0,1]$ and $g_{k}$ are given in Lemma A.2. For $k=1$ this corresponds to $\text{Cov}_{\tilde{\eta}_{\tau}}\left(g(\eta_{1}),\left|\nabla g(\eta_{1})\right|^{2}\right)>0$ suggesting that conditions under which the W-FR could lead to a speed-up lead to a decrease in convergence speed for the FR-W split. This is directly analogous to the Gaussian case (see Proposition 3.2).
6 Conclusion & Future Work
Operator splitting methods lie at the heart of many gradient flow based algorithms, in particular, for discretisations of the Wasserstein–Fisher–Rao gradient flow. We examine errors introduced by splitting schemes whereby the Wasserstein and Fisher–Rao operators are solved separately (and exactly). Such schemes preserve the invariant measure, but produce dynamics that differ from the exact WFR flow. We develop variational formulae for the split schemes and identify a perturbation which modifies the dynamics. It is demonstrated that the order in which these operators are evaluated can have a considerable impact on convergence speed, without affecting $\pi$-invariance. This impact is unsurprisingly more pronounced when the chosen step size is large, but not so large that it is nearly equivalent to evaluating either the Wasserstein or Fisher-Rao flow alone. In practical terms, this means that a carefully chosen splitting can exhibit faster convergence to the target than the exact WFR dynamics and thus a wise choice of splitting can reduce time to convergence without increasing the computational cost. Our work thus highlights that in the context of gradient flow PDEs for sampling one should focus on numerically approximating the split dynamics rather than the exact flow.
We characterise this behaviour in the Gaussian case, where we analyse the evolution of the $\mathrm{KL}$ to $\pi$ using analytic formulas rather than bounds (Section 3). As a generalisation, we consider the strongly log-concave case and show that the exact Wasserstein–Fisher–Rao flow preserves log-concavity uniformly in time, under some conditions (Section 4). This is in sharp contrast with the Wasserstein flow which is known to preserve log-concavity only in the Gaussian case [KOL01].
An upper bound on $\mathrm{KL}$ decay of the exact WFR flow is then characterised in terms of the symmetrised $\mathrm{KL}$, using the derived log-concavity constants and new functional inequalities for Fisher–Rao flows in [CCH+24]. Proposition 5.1 establishes that the WFR flow does in fact reduce the symmetrised KL at a rate which is given by the sum of the rate of the W flow and the rate of the FR flow as conjectured in [DP23]. An analogous result is obtained for the split dynamics where the W flow is evaluated first, then FR, using the variational formula derived in Section 2.
Under appropriate conditions, we obtain a sharper rate of decay of the symmetrised KL for the W-FR flow. This result suggests that operator splitting schemes lead to a speed-up in convergence to the target outside the Gaussian case too. We point out however, that Proposition 5.2 is not sufficient to ensure a speed-up as it only provides an upper bound on the symmetrised KL and not a lower bound, thus preventing us from obtaining an ordering type inequality for the symmetrised KL decay under the two schemes. Nevertheless, results of the form of Proposition 5.2 provide insights on sampling strategies: if the initial distribution $\mu_{0}$ does not satisfy Assumption 4 for $\tau=0$ then it is unlikely that the W-FR split will lead to a speed-up and the FR-W split should be considered instead.
Our analysis assumes that the FR and W operator are evaluated exactly. Due to the large amount of numerical schemes approximating the W flow [DMM19, RS02, SKL20] and the FR flow [LLN19, LSW23, CP25, MM24], we leave the investigation of numerical schemes approximating the split dynamics to future work, and in particular, the study of how numerical errors propagate through the split dynamics.
Finally, we have focused on the impact of operator ordering to improve convergence speed. Further work would also examine how the split step size $\gamma$ could be chosen to maximise this improvement, as e.g. Figure 3.1 demonstrates that the step size should be chosen not too large, but also not too small that the dynamics are close to the exact WFR. Furthermore, adaptive strategies for the numerical approximation of the W-FR flow could be borrowed from the sequential Monte Carlo literature on adaptive tempering [JSD+11, SBC+24] and more recent results on adaptive stepsizes for the W flow [SN25] in investigating optimal step sizes.
Acknowledgements
The authors would like to thank Josh Bon, Sam Power, Florian Maire, Daniel Paulin, Upanshu Sharma for helpful discussions and in particular Sinho Chewi for pointing to the reference [KOL01] and Andre Wibisono for directing us to Lemma 16 in [LMW25]. The authors gratefully acknowledge the mathematical research institute MATRIX in Australia where part of this research was performed.
Appendix A Proofs of variational formulae
A.1 Proof of Proposition 2.2
Here we use the notation $\rho_{1}$ to refer to the unnormalised form of $\nu_{1}$. Denote by $f_{\textrm{G}}$ the unnormalised FR operator,
$f_{\textrm{G}}(\mu):=\mu\log\frac{\pi}{\mu}.$
$\rho_{1}=S_{\textrm{G}}(\gamma,S_{\textrm{W}}(\gamma,\mu_{0})),$
where $S_{\textrm{G}}(\gamma,v)$ denotes the abstract solution operator of $\partial_{\gamma}\rho=f_{\textrm{G}}(\rho)$ initialised at $v$ after time $\gamma$ and $S_{\textrm{W}}(\gamma,v)$ denotes the solution operator of $\partial_{\gamma}\rho=f_{\textrm{W}}(\rho)$. Using the chain rule, the first variation of $\rho_{1}$ with respect to $\gamma$ is (see also (3.4) in [DT13])
$\partial_{\gamma}\rho_{1}=f_{\textrm{G}}(\rho_{1})+\partial_{2}S_{\textrm{G}}(\gamma,\hat{v}_{1})f_{\textrm{W}}(\hat{v}_{1}),$ (1.1)
where $\hat{\nu}_{1}:=S_{\textrm{W}}(\gamma,\mu_{0})$ and $\partial_{i}$ denotes the directional derivative w.r.t. the $i$-th argument. Due to [CHH+23, Appendix B.1], the solution operator $S_{\textrm{G}}$ has an explicit form given by $S_{\textrm{G}}(\gamma,v)=\pi^{1-e^{-\gamma}}v^{e^{-\gamma}}$, so that $\partial_{2}S_{\textrm{G}}(\gamma,v)h$, i.e. the directional derivative w.r.t. the second argument $v\in\mathcal{P}_{2}^{ac}$ in the direction $h\in\mathcal{T}$ is easily computed as,
$\partial_{2}S_{\textrm{G}}(\gamma,v)h=\left.\frac{d}{d\epsilon}S_{\textrm{G}}(\gamma,v+\epsilon h)\right|_{\epsilon=0}=e^{-\gamma}\pi^{1-e^{-\gamma}}v^{e^{-\gamma}-1}h=e^{-t}S_{\textrm{G}}(t,v)hv^{-1}.$
Note that $\partial_{2}S_{\textrm{G}}(\gamma,v)h$ is a linear operator and $x\mapsto\partial_{2}S_{\textrm{G}}(\gamma,v)(x)=e^{-t}(\pi(x)/v(x))^{1-e^{-t}}$ is continuous if $\pi,v$ are continuous and $v(x)>0$. Substituting back into (1.1) yields
$\partial_{\gamma}\rho_{1}=f_{\textrm{G}}(\rho_{1})+e^{-\gamma}S_{\textrm{G}}(\gamma,\hat{\nu}_{1})f_{\textrm{W}}(\hat{\nu}_{1})\hat{\nu}_{1}^{-1}.$ (1.2)
Notice that the r.h.s can be written entirely in terms of $\rho_{1}$ since $S_{\textrm{G}}$ is an invertible operator under our assumptions, with inverse given by
Since $\rho_{1}=S_{\textrm{G}}(\gamma,\hat{\nu}_{1})$, we have using the above that,
$\hat{\nu}_{1}=\pi^{1-e^{\gamma}}\rho_{1}^{e^{\gamma}}.$
Then $f_{\textrm{W}}(\hat{\nu}_{1})$ can be expressed in terms of $\rho_{1}$ since, using shorthand $w:=\nabla\log\frac{\rho_{1}}{\pi}$,
$f_{\textrm{W}}(\hat{\nu}_{1})$ $=\nabla\cdot\left(\hat{\nu}_{1}\nabla\log\frac{\hat{\nu}_{1}}{\pi}\right)$
$=e^{\gamma}(\pi^{-e^{\gamma}+1}\rho_{1}^{e^{\gamma}-1}\nabla\cdot\left(\rho_{1}w\right)+\pi^{-e^{\gamma}+1}\rho_{1}w\cdot\nabla(\rho_{1}^{e^{\gamma}-1})+\rho_{1}^{e^{\gamma}}w\cdot\nabla(\pi^{-e^{\gamma}+1}))$
$=e^{\gamma}\pi^{-e^{\gamma}+1}\rho_{1}^{e^{\gamma}-1}f_{\textrm{W}}(\rho_{1})+e^{\gamma}(\pi^{-e^{\gamma}+1}\rho_{1}w\cdot\nabla(\rho_{1}^{e^{\gamma}-1})+\rho_{1}^{e^{\gamma}}w\cdot\nabla(\pi^{-e^{\gamma}+1}))$
$=e^{\gamma}\pi^{-e^{\gamma}+1}\rho_{1}^{e^{\gamma}-1}(f_{\textrm{W}}(\rho_{1})+\underbrace{\rho_{1}^{2-e^{\gamma}}w\cdot\nabla(\rho_{1}^{e^{\gamma}-1})+\rho_{1}\pi^{e^{\gamma}-1}w\cdot\nabla(\pi^{-e^{\gamma}+1})}_{=:R}),$
and notice that $R$ can be simplified further as
$R$ $=\rho_{1}\left(\rho_{1}^{1-e^{\gamma}}w\cdot\nabla(\rho_{1}^{e^{\gamma}-1})+\pi^{e^{\gamma}-1}w\cdot\nabla(\pi^{-e^{\gamma}+1})\right)$
$=\rho_{1}\left((e^{\gamma}-1)\rho_{1}^{-1}w\cdot\nabla\rho_{1}-w\cdot(e^{\gamma}-1)\pi^{-1}\nabla\pi+\nabla\cdot w\right)$
$=(e^{\gamma}-1)\rho_{1}|w|^{2},$
so that
$f_{\textrm{W}}(\hat{\nu}_{1})=e^{\gamma}\pi^{-e^{\gamma}+1}\rho_{1}^{e^{\gamma}-1}(f_{\textrm{W}}(\rho_{1})+(e^{\gamma}-1)\rho_{1}|w|^{2}).$
Substituting all the derived expressions back into (1.2) yields
$\partial_{\gamma}\rho_{1}$ $=f_{\textrm{G}}(\rho_{1})+e^{-\gamma}e^{\gamma}\pi^{-e^{\gamma}+1}\rho_{1}^{e^{\gamma}-1}(f_{\textrm{W}}(\rho_{1})+(e^{\gamma}-1)\rho_{1}|w|^{2})\pi^{e^{\gamma}-1}\cdot\rho_{1}^{1-e^{\gamma}}$
$=f_{\textrm{W}}(\rho_{1})+f_{\textrm{G}}(\rho_{1})+(e^{\gamma}-1)\rho_{1}|w|^{2}.$
Then by an application of [CP25, Lemma A.1] to obtain the PDE for the normalised density $\nu_{1}$, the claim holds.
A.2 Proof of Proposition 2.3
By similar reasoning as in Section 2.1, the sequential split solution at time $\gamma$ in the order FR-W is given by $\eta_{1}=S_{\textrm{W}}(\gamma,S_{\textrm{FR}}(\gamma,\mu_{0}))$ and its first variation w.r.t. $\gamma$ can be computed as
$\partial_{\gamma}\eta_{1}=f_{\textrm{W}}(\eta_{1})+\partial_{2}S_{\textrm{W}}(\gamma,u_{\gamma})f_{\textrm{FR}}(u_{\gamma}),$ (1.3)
where $\hat{\eta}_{1}:=S_{\textrm{FR}}(\gamma,\mu_{0})$. By Lemma 1 in [DT13] and recalling that $S_{\textrm{W}}$ is a linear operator for which $\partial_{2}S_{\textrm{W}}(t,v)h=S_{\textrm{W}}(t,h)$, we have
$\partial_{2}S_{\textrm{W}}(\gamma,\hat{\eta}_{\gamma})f_{\textrm{FR}}(\hat{\eta}_{\gamma})$ $=f_{\textrm{FR}}(S_{\textrm{W}}(\gamma,\hat{\eta}_{\gamma}))+\int_{0}^{\gamma}\partial_{2}S_{\textrm{W}}(\gamma-\tau,\hat{\eta}_{\gamma})[f_{\textrm{W}},f_{\textrm{FR}}](S_{\textrm{W}}(\tau,\hat{\eta}_{\gamma}))d\tau$
$=f_{\textrm{FR}}(\eta_{1})+\int_{0}^{\gamma}S_{\textrm{W}}\left(\gamma-\tau,[f_{\textrm{W}},f_{\textrm{FR}}](S_{\textrm{W}}(\tau,\hat{\eta}_{\gamma}))\right)d\tau,$ (1.4)
since $S_{\textrm{W}}$ is a linear operator and where $f_{\textrm{W}},f_{\textrm{FR}}](v)$ refers to the Lie commutator (see (3.2a) in [DT13]),
where $D_{\textrm{FR}}$ is the Lie derivative operator associated with $f_{\textrm{FR}}$, i.e. $D_{\textrm{FR}}H(v)=H^{\prime}(v)f_{\textrm{FR}}(v)$ for any unbounded nonlinear operator $H:X\rightarrow\mathcal{T}$, $X\subset\mathcal{P}_{2}^{ac}$ and $H^{\prime}(v)$ denotes the directional derivative at $v$. With a slight abuse of notation, we use $f_{\textrm{W}}^{\prime}f_{\textrm{FR}}(v)$ in place of $f_{\textrm{W}}^{\prime}(v)f_{\textrm{FR}}(v)$ to denote the directional derivative of $f_{\textrm{W}}$ at $v$ in the direction $f_{\textrm{FR}}(v)$. Substituting (1.4) into (1.3) yields that the variation of $\eta_{1}$ takes the form of the usual WFR PDE plus a perturbation term,
$\partial_{\gamma}\eta_{1}=f_{\textrm{W}}(\eta_{1})+f_{\textrm{FR}}(\eta_{1})+\int_{0}^{\gamma}S_{\textrm{W}}\left(\gamma-\tau,[f_{\textrm{W}},f_{\textrm{FR}}](S_{\textrm{W}}(\tau,\hat{\eta}_{\gamma}))\right)d\tau,$ (1.7)
where $f_{\textrm{W}},f_{\textrm{FR}}]S_{\textrm{W}}(\tau,\hat{\eta}_{\gamma})\in\mathcal{T}$. Substituting in (1.13) developed in Lemma A.2 for the Lie commutator leads to (2.11).
$\int_{0}^{\gamma}S_{\textrm{W}}\left(\gamma-\tau,[f_{\textrm{W}},f_{\textrm{FR}}](S_{\textrm{W}}(\tau,\hat{\eta}_{\gamma}))\right)d\tau=\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}\mathcal{G}^{k}(\eta_{1}).$
The result then follows using Lemma A.2.
A.3 Auxiliary Lemmas
Lemma A.1.
Let Assumption 1 hold. Recall $\eta_{1}(x;\gamma)$ denotes the solution of (2.12) over a single step of size $\gamma$. Recall the $k$th commutator $\mathcal{G}^{k}\eta_{1}:=[f_{\textrm{W}},\mathcal{G}^{k-1}\eta_{1}]$ for $k=1,2,\dots$ with $\mathcal{G}^{0}:=f_{\textrm{FR}}$ and $\mathcal{G}^{1}$ as defined in (1.5). The following identity holds
$\int_{0}^{\gamma}e^{(\gamma-\tau)f_{\textrm{W}}}\mathcal{G}^{1}\eta_{1}(x;\tau)d\tau=\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}\mathcal{G}^{k}(\eta_{1}(x;\gamma)).$